### v1

# InstrAug: Automatic Instruction Augmentation for Multimodal Instruction Fine-tuning

Abstract
Fine-tuning large language models (LLMs) on multi-task instruction-following data has been proven to be a powerful learning paradigm for improving their zero-shot capabilities on new tasks. Recent works about high-quality instruction-following data generation and selection require amounts of human labor to conceive model-understandable instructions for the given tasks and carefully filter the LLM-generated data. In this work, we introduce an automatic instruction augmentation method named InstrAug in multimodal tasks. It starts from a handful of basic and straightforward meta instructions but can expand an instruction-following dataset by 30 times. Results on two popular multimodal instruction-following benchmarks MultiInstruct and InstructBLIP show that InstrAug can significantly improve the alignment of multimodal large language models (MLLMs) across 12 multimodal tasks, which is even equivalent to the benefits of scaling up training data multiple times. The code and datasets are available at https://github.com/declare-lab/InstrAug.
1 Introduction
In recent years we have witnessed a surge in the family of large language models (LLMs). Thanks to the scaling factor regarding both data and model size Chowdhery et al. (2023), a series of works have established powerful foundation models which show superior zero-shot performance on the downstream tasks Brown et al. (2020); Touvron et al. (2023a, b); Achiam et al. (2023); Team et al. (2023); Bai et al. (2023). Instruction fine-tuning (IFT), which aims to “teach" models to follow natural language instructions, has been proven to be an effective learning paradigm to further enhance LLMs’ generalizability Sanh et al. (2021); Wei et al. (2021); Chung et al. (2022); Taori et al. (2023).
When deploying the instruction-tuned LLMs on unseen tasks, users from various backgrounds may write instructions with diverse wording styles to test the performance of the same task. In light of this, high-quality instruction-following data covering a wide range of wording styles are required in the alignment of large language models. The conventional approach to obtain a large instruction-following dataset resembles data annotation, which recruits crowdsourced workers who engage in NLP research and possess rich knowledge about the target tasks to write instructions for each task/instance Sanh et al. (2021); Wang et al. (2022b). Zhou et al. (2023) further eliminates redundant annotations, proving that IFT on only well-crafted 1,000 instructions can align as huge as 65B models with user preferences. While achieving a comparable level to the state-of-the-art baselines like GPT-4 Achiam et al. (2023), these LLMs instruction-tuned by hand-written data inevitably necessitate numerous iterations of thoughtful consideration and checking, which cost substantial human labor. Although recent works have explored automatic instruction generation through in-context learning Wang et al. (2023b); Honovich et al. (2023); Taori et al. (2023), the generated instruction-data pairs still require meticulous identification and filtering by experienced workers to yield a practical dataset.
The same issue exists in multimodal tasks. Multimodal IFT (MIFT) has demonstrated remarkable efficacy in addressing visual question answering tasks. Some of the instruction tuning datasets such as MIns+ Xu et al. (2023b) contain more than 500K instances but pretty few instructions. To break through the bottleneck, we argue that Enriching the training data with diverse task-specific instructions can enhance the effectiveness of MLLMs to a range of zero-shot task instructions and propose a completely automatic instruction-following dataset expansion framework InstrAug for MIFT. By harnessing an LLM and starting from straightforward meta-prompts Suzgun and Kalai (2024), we expand the hand-written instruction templates to a size 30 times larger. We fine-tune the two base models on the augmented dataset that incorporates the generated instruction templates. The results on two benchmarks that cover a broad range of task types manifest that InstrAug can facilitate the IFT learning process, particularly in scenarios where the training data is limited. We find augmenting the dataset with about 1500 generated instructions can boost MLLM’s performance by 2-3%. Moreover, it exhibits better adaptability to instructions in different language styles and nearly catches up with the result by expanding the dataset 10 times (Figure 1). To put it in a nutshell, our contributions are as follows:
We present InstrAug, an automatic and labour-free dataset expansion framework for MIFT. InstrAug relies purely on rule-based filters to remove invalid generations and adopts a self-adaptive sampling strategy to balance instructions’ consistency and diversity in the resulting dataset. Compared to previous peer works, InstrAug almost requires no human labor and can be done by a single person.
We expand a raw MIFT dataset MultiInstruct (MIns) to MIns+ (this name means it is an upgraded version of MultiInstruct) using InstrAug. Fine-tuning two MLLMs on MIns+ yields promising results on different task types whose instruction templates cover a broad range of wording styles.
We further conduct a comprehensive empirical study based on the obtained results and propose our insights from multiple aspects.
2 Related Work
2.1 Instruction Finetuning on MLLMs
Finetuning pretrained large language models on human-written multi-task instruction-following data has shown unprecedented power for zero-shot task generalization as well as alignment with human intentions to accomplish real-world tasks Wei et al. (2021); Wang et al. (2023b, 2022b); Ouyang et al. (2022); Longpre et al. (2023); Liu et al. (2023a). As large visual foundation models like Flamingo Alayrac et al. (2022), BLIP Li et al. (2022, 2023c) and SegmentAnything Kirillov et al. (2023) emerge, many visual-language tasks can be formulated in a sequence-to-sequence prediction manner and effectively solved by these models, and consequently, IFT has been successfully extended to the multimodal field and its efficacy has been verified Liu et al. (2023a); Zhang et al. (2023); Dai et al. (2023); Han et al. (2023); Li et al. (2023a). Most of these works employ rule-guided human-written instructions through crowdsource workers, which are not cost-effective if the required instruction size is huge.
2.2 Automatic Instruction Generation
Since instruction curating is costly nowadays, researchers seek automatic data generation methods to cut down the expenses. Unnatural instructions Honovich et al. (2023) and Self-instruct Wang et al. (2023b) firstly exploit GPT-3 (text-davinci) and in-context learning to generate new I/O pairs and corresponding instructions. This method has been applied to produce more advanced instruction-following models like Alpaca Taori et al. (2023) and LLaMA-adapter Zhang et al. (2023). Li et al. (2023d) train a model to generate instructions inversely for the input instances and iteratively selects high-quality data to train this instruction generation model the in next stage. Zhou et al. (2023) prove as few as 1,000 carefully curated instructions are sufficient to yield decent zero-shot performance on downstream tasks. These works inevitably utilize a number of trained human annotators. The other line of works Xu et al. (2023a); Sun et al. (2023); Liu et al. (2023b) proposed to use a few conceived abstract principles or augmentation guidelines to refine existing instructions in order to obtain model-understandable instructions. However, these works concentrate on the augmentation of input text by increasing the difficulty or broadening the question scope instead of task description data, which may not be directly applied to augment task descriptions.
3 Method
3.1 Multimodal Instruction Fine-tuning (MIFT)
MIFT uses instruction-following data in the training set to tune a base model with parameter , and expects the model to perform well on the unseen test tasks. Formally, given the MIFT dataset , where is the dataset collection of tasks in the individual split. to satisfy the zero-shot setting. Each task-specific dataset consists of input-output pairs and an instruction template sets , where means there is usually more than one template for each task. Each instance has been instantiated before being fed to the base model using instruction template :
The tuning procedure resembles the classical sequence-to-sequence generation, where we maximize the likelihood of ground truth sequence of length in an auto-regressive fashion:
The model is then assessed on , where the score is calculated with the task-specific criterion between predicted output and truth tokens and averaged across the dataset .
3.2 Generation Process
The overall pipeline of our proposed framework is illustrated in Figure 2. We utilize LLaMA2-13B-Chat to generate new instructions. At the beginning, we initialize the LLM with meta-prompts, which provide guiding instructions to steer the LLM in generating new instructions. The inputs to the LLMs comprise both the guiding instructions and original instructions, encompassing raw instructions from MIns (solid line) and filtered valid outputs from the preceding step (dashed line, as we consider only valid inputs likely to yield valid outputs). Upon gathering the outputs from the LLM, we apply a sequence of concise rule-based filters to eliminate invalid content and inconsistent instructions. Subsequently, we construct MIns+ using the generated instruction templates and the raw instances. Finally, we fine-tune the base model on MIns+ (not depicted in the figure).
3.2.1 Meta-Prompts
We explore two types of meta-prompts: direct prompting and bootstrapping. Direct prompting involves providing a clear, human-conceived prompt that explicitly communicates the objective of generating new instructions from the given one to the LLM. For instance, an example prompt could be, "Rephrase the given instruction." In contrast, bootstrapping compels the LLM to generate a predetermined number (e.g., 10 in our experiments) of instructions that can guide the LLM in generating direct prompting instructions. The specific input-output examples can be found in Section B.1.
For both types of meta-prompts, we adhere to the common practice Touvron et al. (2023b) of prepending a system prompt such as "You are a helpful assistant." (Figure 2) to the generation prompt. The generation procedure can iterate and repeat numerous times, where the valid output from the last step serves as the input in the new iteration. As the loop proceeds, the occurrence of duplicated outputs increases. Consequently, we halt the process after several iterations once we have amassed the desired number of instructions. Detailed prompt template formatting and response examples are available in the appendix.
In addition to iterative generation, inspired by Wang et al. (2023b), we also experiment with sampling from different temperatures to further diversify the instruction set. In subsequent discussions, we refer to this variant as , while the iterative collection method is denoted as .
3.2.2 Handling Placeholders
One challenge during generation is how to deal with instructions with placeholders—blanks to be filled during instantiating time. Such placeholders can be simply expressed as format string variables in Python, but can not be properly handled by an LLM. For example, one of the prompts for the Grounded Captioning task is “Describe the content of {region_split_token.join(region)} in the picture.”, where “region_split_token” and “region” are item names from an instance and can be initialized as the format string in Python when loading the instance. If we directly ask the model to rewrite, the output could be “Give a description for {list of region}”, since the LLM may treat the content within the brackets as regular words/phrases and rewrite them as well. This outcome incurs value error when built with examples as format strings in the code.
To alleviate this issue, we devise a Placeholder-Protected Generation (PPG) algorithm that combines the technique of masking and controlled generation, as illustrated in Algorithm 1. We substitute placeholders with a predefined list of “atomic” (i.e., hardly to be modified in generation) masks before feeding them into the language model, then we map them back to these placeholders in the generation results. Specifically, we found single alphabet has promising atomic properties. Hence we choose these simple elements as protectors. During generation time, we append the restriction command “Keep the content within brackets (including ‘{}’) unchanged.” to the generation instruction. By combining these two add-ons, the model can produce valid augmentations in most cases.
3.2.3 Postprocessing
We apply the following rule-based filters to the primary results: 1) Duplication examination and removal through perfect match. 2) For original instructions containing placeholders, we preserve those in an unordered match with the original ones. 3) Length-based filters to overcome non-aligned responses since we found these responses are much longer than the normal ones (See Section B.3 for examples). We find this hallucination removal filter is extremely effective when the instruction set is large (See our ablation study in Section 4.5). Finally, we gather all valid instructions for the building of the augmented dataset. The details of the result dataset can be found in Appendix C.
3.3 Fine-tuning on MIns+
3.3.1 Source Dataset
We curate MIns+ by sampling from the training dataset collection of MultiInstruct Xu et al. (2023b), which consists of 62 tasks (53 training tasks and 9 evaluation tasks) based on 21 existing datasets. We use the same split of training and testing tasks as in MultiInstruct. Since it is unknown how many instances exactly sampled from each task, we respectively sample two versions in MIns+—1K and 10K instances per task (59K and 564K in total), which also provides an opportunity to study our method’s effectiveness under different dataset sizes. Unless specifically mentioned, each instance only appears once in the final dataset.
3.3.2 Dataset Construction
The next step is to amplify the raw dataset with generated instructions. A straightforward practice is to randomly pick one instruction template (with repetition) and one instance (without repetition) and combine them to form the new dataset. However, from the perspective of data augmentation, this model-agnostic method neglects the distribution of these instruction templates in the MLLM’s textual encoder’s hidden space. Therefore, we develop an adaptive sampling procedure within each task’s scope. Concretely, we first assign probability and to the original and generated instruction set and , respectively, where is a preset parameter and is the cardinality of set . For original instructions, we deem they contribute equally to the training so that the probability to pick each template from for task is . For generated instructions, we calculate for each of them a score , and the probability is over a softmax normalization. By default in our experiments to ensure a balanced mapping.
To compute , a formulation should balance well between consistency and diversity Sanh et al. (2021). Consistency means that the generated instruction can convey precise task description as its origin. Diversity denotes that the training procedure has a comprehensive coverage over the instruction wording space so that for various instruction templates the model can produce similar results. Combining these two factors, we use the following expression to score each generated instruction:
where is the similarity function, is the embedding of produced from the MLLM’s text encoder. The first item quantifies the similarity between the generated instruction and its origin (consistency), while the second item measures the divergence (negative similarity, or diversity) between the instruction and randomly sampled counterparts generated from the same .
4 Experiments
4.1 Baselines and Base Models
We conduct experiments on the following two base models and evaluation tasks after finetuning on different versions of MIns+. Specifically, we compare the results with a) zero-shot performance (the checkpoints of base models at step 0); b) MIns finetuned base models; c) InstructBLIP finetuned on the 15M tuning set Dai et al. (2023), denoted as .
OFA
Wang et al. (2022a) is a task-agnostic sequence-to-sequence learning framework, which supports both unimodal and multimodal tasks. It is pretrained on 20M publicly available image-text pairs but can show great transferability to unseen tasks and domains. In our experiments, following Xu et al. (2023b), we finetune OFA-large which contains 870M parameters during evaluations.
InstructBLIP (Vicuna-7B)
Dai et al. (2023) is one of the earliest trial on MIFT. It is built on pretrained BLIP-2 Li et al. (2023c) and Vicuna Chiang et al. (2023), a finetuned version of LLaMA on user-shared conversations. InstructBLIP is tuned on a sampled subset of 15M instances covering 7 major tasks and 13 datasets. During the fine-tuning stage, only the fully connected layer which projects the instruction-aware image features into Vicuna’s input space, is updated, while both the image-encoder and Query Transformer (Q-Former) in BLIP-2 and the upside LLM (Vicuna-7B) are frozen.
4.2 Tasks and Metrics
We select 12 tasks from the test set of two representative MIFT benchmarks, MultiInstruct Xu et al. (2023b) (9 tasks) and InstructBLIP Dai et al. (2023) (3 tasks) for evaluation, as shown in Table 10. Among these tasks, VSR Liu et al. (2023a) and Visdial Das et al. (2017) are included in both benchmarks (counted as MultiInstruct tasks). We test for extant templates on MIns tasks, which can be found in Xu et al. (2023b). For IBLIP-Bench multiple-choice tasks (ScienceQA and IconQA), we reconstruct another template in MIns style for these tasks based on the templates from the IBLIP paper Dai et al. (2023). The detailed templates can be found in Section D.5.
4.3 Main Results
We report the results of MIns tasks on OFA-Large and InstructBLIP in Table 1 and Table 2. There are several key findings from the two tables:
First, finetuning both models on MIns and MIns+ acquires significant gain on the held-out tasks over base models which only experience the pretraining stage, demonstrating that MIFT can enhance the model’s task generalisability.
Second, MIns+ outperforms MIns on almost all tasks and settings (except VisDial on OFA-564K), which shows that tuning on datasets composed of both raw and augmented instructions casts a positive effect on the completion of downstream tasks. A surprising discovery is that the result of MIns+ (59K) is competitive with (42.45 vs. 43.89, 1.44 behind OFA-Large) and even surpasses (37.33 vs. 36.87, 0.46 points ahead of IBLIP) that of MIns (564K), implying that the tuning with augmented instruction set can achieve the similar level of performance as dataset expansion (approximate 10 times). Nevertheless, this conclusion is in terms of the overall score, at the task level some performance gaps cannot be bridged through merely instruction augmentation (e.g., GVQA on OFA-Large and CVQA on both models).
Lastly, using the same instruction set, although performance on most of the tasks can be improved from the 59K tuning set to the 564K counterparts, there are some “abnormal” cases that the performance tends to be worse. This effect depends on the target model. For instance, the accuracy of DC reduces by 0.23.9 points on OFA but continues to rise on IBLIP. Besides, under the same setting, the performance improvements vary across tasks. We will have a further discussion on this topic in Section 4.4 and Section 5.1.
Size | Method | VTE | TVQA | VisDial | CVQA | DC | GVQA | VE | VSR | NLVR | Avg. |
Rouge-L | Acc. | ||||||||||
0 | OFA-Large | 14.97 | 10.79 | 29.75 | 0.64 | 4.96 | 0.00 | 43.65 | 52.03 | 53.98 | 23.42 |
59K | +MIns | 29.44 | 20.75 | 35.14 | 23.24 | 49.28 | 20.48 | 54.22 | 56.27 | 62.60 | 39.05 |
+MIns+ | 35.33 | 21.52 | 35.55 | 22.70 | 54.94 | 20.63 | 54.24 | 67.48 | 69.64 | 42.45 (+3.40) | |
+ | 30.11 | 20.83 | 34.32 | 23.16 | 51.94 | 20.62 | 54.12 | 67.59 | 71.77 | 41.61 (+2.56) | |
+ | 31.50 | 20.74 | 33.30 | 23.28 | 52.44 | 20.38 | 54.48 | 67.83 | 69.36 | 41.48 (+2.43) | |
564K | +MIns | 41.33 | 24.49 | 36.18 | 26.84 | 47.68 | 49.25 | 50.74 | 63.96 | 54.52 | 43.89 |
+MIns+ | 45.43 | 24.62 | 34.98 | 29.16 | 51.06 | 49.64 | 52.76 | 66.74 | 54.70 | 44.92 (+1.03) | |
+ | 45.63 | 23.36 | 34.30 | 30.02 | 53.84 | 53.73 | 52.18 | 66.11 | 54.34 | 45.95 (+2.06) | |
+ | 51.71 | 24.55 | 35.08 | 27.52 | 52.24 | 56.12 | 53.06 | 66.72 | 55.74 | 46.78 (+2.89) | |
0 | IBLIP | - | 3.74 | - | 24.26 | 15.80 | 22.70 | 33.42 | 49.82 | 50.10 | 28.55 |
59K | +MIns | - | 17.14 | - | 27.46 | 17.44 | 24.00 | 51.76 | 53.60 | 54.12 | 35.07 |
+MIns+ | - | 17.84 | - | 27.54 | 22.46 | 25.30 | 53.30 | 56.08 | 56.28 | 36.97 (+1.90) | |
+ | - | 18.01 | - | 28.10 | 21.10 | 24.14 | 53.64 | 54.74 | 56.48 | 36.60 (+1.53) | |
+ | - | 18.29 | - | 27.62 | 22.72 | 25.36 | 54.84 | 55.66 | 56.80 | 37.33 (+2.26) | |
564K | +MIns | - | 15.12 | - | 27.66 | 29.28 | 24.20 | 50.44 | 52.84 | 56.26 | 36.87 |
+MIns+ | - | 17.19 | - | 29.24 | 34.82 | 25.48 | 54.56 | 57.16 | 57.30 | 39.39 (+2.52) | |
+ | - | 16.67 | - | 27.90 | 33.30 | 26.36 | 53.90 | 56.06 | 56.81 | 38.71 (+1.84) | |
+ | - | 16.09 | - | 27.88 | 36.48 | 25.86 | 55.34 | 56.52 | 57.22 | 39.34 (+2.47) | |
15M | + | - | 22.26 | - | 25.40 | 26.72 | 23.86 | 35.18 | 52.04 | 52.22 | 33.95 |
Size | Method | SciQA | IconQA | VizWiz | Avg. |
0 | OFA-Large | 3.90 | 16.92 | 7.76 | 12.42 |
564K | +MIns | 26.18 | 38.70 | 18.67 | 27.85 |
+MIns+ | 27.32 | 38.25 | 19.88 | 28.48 (+0.63) | |
+ | 26.69 | 38.45 | 18.94 | 28.03 (+0.18) | |
+ | 26.41 | 39.40 | 19.39 | 28.40 (+0.55) | |
0 | IBLIP | 38.29 | 37.71 | 12.19 | 29.40 |
564K | +MIns | 49.03 | 39.85 | 24.73 | 37.87 |
+MIns+ | 49.75 | 40.22 | 27.87 | 39.28 (+1.41) | |
+ | 49.26 | 39.91 | 27.96 | 39.04 (+1.17) | |
+ | 49.12 | 40.17 | 25.92 | 38.40 (+0.53) |
4.4 Scaling Factors
We also notice the scaling factors from the perspective of the dataset (including instruction and instance) and model size.
Dataset
As shown in Table 1, the performance can be enhanced by expanding the size of either instruction set (MIns to MIns+) or instance set (59K to 564K) for tuning. Under the same instance set size, using a larger instruction set can boost the overall performance (MIns to MIns+). Nevertheless, the marginal gain gradually decreases and can even be reverted to negative. For example, the marginal gain on OFA 59K, from MIns to MIns+ is 3.40, but is and from MIns+ to and , respectively. In contrast, under the same instruction set (MIns+ and ), MLLM’s scores significantly elevate in a larger instance set. This outcome perhaps stems from more precise human-annotated datasets compared to automatically generated instructions that roughly match the original meaning, implying that data are more essential and instructions are auxiliary ingredients for an IFT training set.
Model
The scaling factor similarly exists in terms of MLLM size. Leveraging the 59K instance set, the 870M OFA earns even higher improvement than 7B+ IBLIP on both raw MIns and instruction-augmented (MIns+ series) tuning set, indicating that a tiny dataset composed of diverse instructions is sufficient for smaller models to achieve decent performance, while larger models (IBLIP) requires a larger dataset (564K) to fulfill similar improvement.
Setting | MIns+ | ||
OFA-Large (59K) | 42.45 | 41.61 | 41.48 |
+Ordered Match | 41.87(-0.58) | 41.24(-0.37) | 41.05(-0.43) |
-Length Filter | 42.38(-0.07) | 40.85(-0.76) | 40.67(-0.81) |
OFA-Large (564K) | 44.92 | 45.95 | 46.78 |
+Ordered Match | 44.21(-0.71) | 45.16(-0.79) | 45.81(-0.97) |
-Length Filter | 44.77(-0.15) | 44.99(-0.96) | 45.61(-1.17) |
4.5 Ablation Study
We consider the following ablative settings regarding the post-processing filters: 1) For placeholder examination, we substitute the unordered matching with ordered matching. That is, if placeholders of the raw instruction are in the order “”, then only the outputs organized in the same order are preserved, and those which look like “” and “” will be eliminated. 2) We remove the length filter to allow hallucinated output to participate in the construction of MIns+. As illustrated in Table 3, forcing ordered matching or discarding the length filter downgrades the overall performance, which implies that shuffling placeholders is an effective augmentation for the model, and the length filter removes harmful hallucinated outputs. Moreover, we note that the extent of negative impact is dependent on the augmented instruction set size. The drop is more severe when enforcing ordered match in MIns+, but on and is larger when disabling the hallucination filter.
5 Further Analysis
In this section, we first seek some task attributes as potential factors to impact the performance of each task on MIns+. Then we explore how the sampling strategy in Section 3.3.2 can affect the overall performance on evaluation benchmarks.
5.1 Potential Impact Factors to MIns+
As mentioned above, InstrAug can improve the performance on the benchmark level. Still, we are curious about what dominates and how much each individual task can benefit from instruction augmentation. To probe the potential factors at the task level, we define the following task particulars:
direct_question (bool): if the instruction template directly tells the model what it should respond to. For example, a template of image captioning “Describe what you see in the image” proposes the entire question to the task. Usually, this type of template does not contain any instance-specific placeholders.
option_inclusive (bool): if the template provides a placeholder for options. This mainly appears in multi-choice question-answering tasks.
template_text_proportion (float in ): the dataset-wise average proportion of template text length (excluding placeholders) in the entire instantiated instruction.
We display the average gain for each task on OFA (59K) together with the individual attributes in Figure 3. The Pearson Correlation Coefficient between instance text proportion and the average performance gain is 0.639, which uncovers a strong correlation between the relative length of instance text to the extra gain by InstrAug. Moreover, tasks with shorter instance text fall behind those with longer instance text. A probably reasonable explanation for this is that MLLMs attend excessively to the task-irrelevant templates, which could perturb the gradient to an inappropriate direction and scale Von Oswald et al. (2023).
When the problem to solve has been entirely enclosed in the template like VTE, the task can benefit more from InstrAug, since for such tasks in the training set InstrAug is imposed directly on the question rather than modifier text. Compared to those whose instructions contain options after instantiation, tasks without visible options (TextVQA and VisDial) are susceptible to deterioration on MIns+, possibly because instruction augmentation can help consolidate the copy-from-input ability by modifying the surrounding context, while pure generative tasks do not rely on this ability to predict answers.
5.2 Sampling Strategy
We investigate the impact of the sampling strategy proposed in Section 3.3.2, mainly study on the hyperparameter . There are 4 different values evaluated: , where is the default task-specific value in our implementations. Evaluations are conducted on MIns+ since the instruction set is sufficiently small to guarantee that and deviate from distantly). According to Figure 4, when sampling with a larger or smaller , the average score on evaluated tasks decreases apparently. The performance becomes even worse when using a fixed large value of for all tasks (0.5). These records reflect the importance of the balance between diversity and consistency in sampling, and our devised simple strategy reaches this goal.
6 Conclusion
In this paper, we propose InstrAug, a completely automatic instruction augmentation framework for large instruction-following datasets with very scarce hand-written instructions. The framework purely relies on an arbitrary LLM and simple meta-prompts. We specially devise the generation process and rule-based filters based on the characteristics of raw instructions and model outputs. Experimental results on two MLLMs across a broad range of tasks demonstrate our proposed method can significantly enhance the overall performance compared to baselines using non-augmented datasets, which also shows the improved robustness of human-written instructions in different wording styles.
Limitations
In InstrAug, we apply a series of rule-based filters to remove invalid outputs that could do evil to the dataset construction and tuning process, as well as a heuristic sampling strategy to balance the distribution of generated instructions in different versions of MIns+. However, InstrAug works on augmentation with less focus on other dataset surgery techniques such as data selection Xia et al. (2024) and dataset pruning Yang et al. (2022), which we believe may further improve the current performance by combining them with InstrAug. We hold this thought as our future research direction.
Ethical Considerations
We clarify ethical considerations in two aspects. First, InstrAug is a fully automatic augmentation framework and no crowdworker is hired. Second, regarding the generated process from LLM (LLaMA-2-13b-Chat), since it is an open-sourced model with safety-tuning, and our instructions do not contain question or instructions that deliberately elicit harmful output from the LLM. We have thoroughly checked the augmented instruction set—apart from possible hallucinations, most of the output follows the guiding instruction and just modifies the extant instructions in MIns. Moreover, we provide reference inputs and outputs in the appendix to make the generation process transparent and public, and commit to release the instruction set for public inspection.
Appendix A More Experimental Results
We also test IBLIP (594K MIns versions) on the Massive Multi-discipline Multimodal Understanding and reasoning (MMMU) benchmark Yue et al. (2023), as shown in Table 4. We can see that IBLIP tuned on augmented MIns surpassses non-augmented baselines on both validation and test sets. Notably, MMMU uses similar template style as IBLIP-bench for evaluation Yue et al. (2023), but the gain on MMMU is more significant than IBLIP-bench (Table 2), which provides a supplementary evidence that pure template augmentation has the chance to benefit tasks with very concise templates.
Business | Science |
(900) | (10,500) | (1,163) | (1,428) | (2,426) | (1,752) | (947) | (2,784) | |||||||||||||
OpenFlamingo2-9B Awadalla et al. (2023) | 28.7 | 26.3 | 31.7 | 23.5 | 26.3 | 26.3 | 27.9 | 25.1 | ||||||||||||
Kosmos2 Peng et al. (2023) | 24.4 | 26.6 | 28.8 | 23.7 | 26.6 | 27.2 | 26.3 | 26.8 | ||||||||||||
Fuyu-8B Bavishi et al. (2023) | 27.9 | 27.4 | 29.9 | 27.0 | 25.6 | 27.0 | 32.5 | 26.4 | ||||||||||||
MiniGPT4-Vicuna-13B Zhu et al. (2023) | 26.8 | 27.6 | 30.2 | 27.0 | 26.2 | 26.9 | 30.9 | 27.2 | ||||||||||||
LLaMA-Adapter2-7B Zhang et al. (2023) | 29.8 | 27.7 | 35.2 | 25.4 | 25.6 | 30.0 | 29.1 | 25.7 | ||||||||||||
Otter Li et al. (2023b) | 32.2 | 29.1 | 37.4 | 24.0 | 24.1 | 29.6 | 35.9 | 30.2 | ||||||||||||
CogVLM Wang et al. (2023a) | 32.1 | 30.1 | 38.0 | 25.6 | 25.1 | 31.2 | 41.5 | 28.9 | ||||||||||||
InstructBLIP-T5-XL Dai et al. (2023) | 32.9 | 30.6 | 43.3 | 25.2 | 25.2 | 29.3 | 45.8 | 28.6 | ||||||||||||
BLIP-2 FLAN-T5-XL Li et al. (2023c) | 34.4 | 31.0 | 43.0 | 25.6 | 25.1 | 31.8 | 48.0 | 27.8 | ||||||||||||
Qwen-VL-7B Bai et al. (2023) | 35.9 | 32.9 | 47.7 | 29.8 | 25.6 | 33.6 | 45.3 | 30.2 | ||||||||||||
LLaVA-1.5-13B Liu et al. (2023a) | 36.4 | 33.6 | 49.8 | 28.2 | 25.9 | 34.9 | 54.7 | 28.3 | ||||||||||||
InstructBLIP-T5-XXL Dai et al. (2023) | 35.7 | 33.8 | 48.5 | 30.6 | 27.6 | 33.6 | 49.8 | 29.4 | ||||||||||||
BLIP-2 FLAN-T5-XXL Li et al. (2023c) | 35.4 | 34.0 | 49.2 | 28.6 | 27.3 | 33.7 | 51.5 | 30.4 | ||||||||||||
InstructBLIP-Vicuna7B Dai et al. (2023) | 27.6 | 24.8 | 29.0 | 25.9 | 22.2 | 25.0 | 25.4 | 24.6 | ||||||||||||
+MIns | 27.4 | 25.7 | 29.1 | 24.6 | 22.8 | 25.5 | 31.6 | 25.3 | ||||||||||||
+MIns+ | 28.2 | 26.8 | 34.3 | 23.9 | 23.4 | 25.8 | 32.9 | 26.7 | ||||||||||||
+ | 29.1 | 26.9 | 33.5 | 25.3 | 22.8 | 26.2 | 31.2 | 26.1 | ||||||||||||
+ | 28.5 | 26.5 | 33.1 | 23.2 | 23.5 | 26.1 | 30.6 | 26.5 |
Appendix B Generation Details
B.1 Guiding Instruction Generation
We first display the details of the guiding instruction generation process, during which we query our LLM with a simple meta-prompt, as shown in Table 5. We omit the irrelevant output in front of and behind the 10 itemized guiding instructions.
B.2 New Instruction Generation
To leave an intuitive impression on the new instruction generation process. The generated instructions (“Gen”) are produced by feeding the corresponding guiding instruction from Table 5 along with the raw instruction (“Raw”) into the LLM. Formally, we use the simple template in Table 6:
where “Guiding Instruction” is 1 of the 10 instructions that we generate from the first step, “Constrain Instruction” only appears when the input instructions have placeholders (it is masked through Algorithm 1) to enforce the model preserve highlighted masks, “[TEXT]” is a special token that works as the delimiter to help the model recognize the boundary of input text, “Raw Instruction” is the raw instruction to be modified.
B.3 The Intuition behind Length Filter
In the generation process, we found there are many cases that which LLM produces hallucinated output. For example, when asking the model to rewrite the following instruction:
Generate some text to describe the image.
Instead of generating a plausible paraphrasing, the output could be like this:
The photograph depicts a majestic landscape with a serene lake in the foreground, surrounded by towering trees with vibrant foliage. The water’s edge is adorned with smooth rocks and lush vegetation, creating a picturesque setting. In the distance, the sky is painted with hues of blue and white, adding a touch of elegance to the scene. The overall ambiance is one of tranquility and natural beauty, inviting the viewer to pause and appreciate the splendor of the outdoors.
In this response text, the model describes an image instead response to our inquiry, and we do not present any image here. This is abnormal behaviour called hallucination Ji et al. (2023). The text here is possibly seen by the model in advance during the pretraining or tuning stage. We particularly note that such hallucinated output is always much longer than the original instructions (78 and 7 in this example). Hence we can rule out those extremely long generated texts to eliminate hallucinations.
[System Prompt] |
You are a helpful assistant. |
[User] |
Generate 10 instructions about how to rephrase short text. Your response should be cononical and formatted as enumerations. |
[System] |
Please follow the description of the given instruction to modify the input text behind the token [TEXT].
{Guiding Instruction} {Constrain Instruction} |
[TEXT]: {Raw Instruction} |
Rewriting Ins (ID) | Task | Type | Instruction Text |
Use synonyms (1) | Sentiment Understanding | Raw |
What is the caption of the image? |
Gen |
What is the description of the image? |
Object Region Match | Raw |
Is the object {text} in {regions}? {options} |
Gen |
Is the object {text} located in {regions}? {options} |
Rearranging Phrases (2) | Image Caption | Raw |
Please analyze the sentiment depicted in the image and answer the question. {question} {options} |
Gen |
Please answer the question after analyzing the sentiment depicted in the image. {question} {options} |
Object Region Selection | Raw |
In this task, you are given the coordinates of some rectangular region in the image. You need to first localize each rectangular region and then identify what is the object in the region. The region is {regions}. |
Gen |
The task at hand is to localize and identify objects within a specified rectangular region in an image. The region of interest is {regions}. To complete the task, you will need to first localize the region and then identify the object within it. |
Simplyfy Language (3) | Grounded Captioning | Raw |
What is the content of {regions}? |
Gen |
What is inside {regions}? |
Visual Attribute | Raw |
Decide which option is the attribute of the object in the given region. Region: {regions} {options} |
Gen |
Determine the attribute of the object in the specified region: {regions} {options} |
Remove Redundancy (4) | Image Caption | Raw |
In this task, you will look at the image and briefly describe the image. |
Gen |
Briefly describe the image you see. |
Object Grounding | Raw |
Given the description of an object, generate the region that contains this object. The description is: {text} |
Gen |
Please generate the region that contains the object described by ’{text}’. |
Use Active Voice (5) | Factual Checking | Raw |
Deicide if the claim can be supported by the image and the context. Context: {context} Claim: "{text}" {options} |
Gen |
Evaluate the validity of the claim {text} {options} based on the image and the surrounding context {context} . |
Object Identification | Raw |
In this task, you are required to answer a question about the appearance of an object.{question} {options}. |
Gen |
For this task, you need to describe the appearance of an object. Please provide your answer in the format of {question} {options}. |
Appendix C MIns+ Details
C.1 Statistics of MIns+
Instance pool of MIns+ is same as MultiInstruct, which encompasses 62 tasks in total (53 tasks in training set). Among the 62 tasks, 34 are predefined existing tasks while the remaining 28 are derived from the annotations on the existing tasks, such as object matching from Visual Genome Krishna et al. (2017) and non-overlapping region selection from MSCOCO Lin et al. (2014).
We list the statistic of MIns and MIns+ in Table 8. Here samples under 11 different temperatures (0.5 to 1.0, with an interval of 0.05). only iterates twice at the default temperature 0.6, since the number of valid instructions increases exponentially (see Table 8, MInsMIns+). From the table we can capture a trend of increment in length as the size grows, but this is restricted to within 15% due to the effect of length filters.
Dataset | MIns | MIns+ | ||
#Instr | 329 | 1855 | 9566 | 8724 |
Avg. Len | 14.66 | 15.46 | 16.78 | 16.50 |
C.2 Examples of Rewritten Instructions
We give some examples of the generated new instructions with corresponding rewriting instructions and original text in Table 7. From the table, we can find that the LLM (LLaMA-2-13b-chat) basically follows its self-generated guiding instruction to produce output.
C.3 Length Distribution
We plot the length distribution of instructions from MIns and three MIns+ datasets, as shown in Figure 5. Instructions whose length exceeds 60 accounts for less than 1% in all datasets so we cut off the histogram within the range . We find that even smaller expansion version MIns+ can have a more smooth distribution than MIns, and larger size MIns+ ( and ) continue to flat the length distribution histogram.
C.4 Wording Characteristics
To observe the wording diversity of MIns+ and comparing the shift with MIns, we further plot the distribution of instruction wording—the first two words of highest frequency in Figure 6 and Figure 7. These figures explicitly reveal a shift from raw wording frequency distribution to generated wording frequency distribution, which appears in
Change in the area of sectors which represent the relative frequency in the overall dataset. For example, the area of “In” and “Given” diminish. Some pairs of words even swap the relative positions that imply the frequency ranking (frequency descends counter-clockwisely), such as “The” and “What” which respectively rank 3rd and 5th in MIns , but rank 5th and 3rd in MIns+.
New sentence patterns created in MIns+, such as those which begin with “For”, “You” and “Please”, etc.
Appendix D Training and Evaluation Details
D.1 Hardware and Hyperparameters
We tune OFA on 4 A6000 GPU of 46G RAM and IBLIP on 2 A100 GPU of 80G RAM. We follow most of the default settings in each tuning version except the batch size for gradient accumulation. The hyperparameter setting for both models can be found in Table 9.
Model | OFA | IBLIP |
Learning Rate | 1e-5 | 1e-5 |
Optimizer | Adam | AdamW |
Adam | 0.9, 0.999 | 0.9, 0.999 |
Epoch | 8 | 8 |
Batch Size | 64 | 128 |
Gradient Accumulation | 4 | 8 |
Image Size | 480480 | 224224 |
FP16 | True | True |
D.2 Tuning Details
We fine-tune the targeted models on MIns+ for 5 epochs, utilizing the default hyperparameters settings. For the 1K/10K (per task) version, the total instance pool size is 59K/564K (with some tasks having fewer than 10K samples). We validate the checkpoint of each epoch for the 59K version and every 2K steps for the 564K version.
D.3 More Details of Evaluation Tasks
Generally, the tasks can be categorized into three types: 1) Generative task, including Visual Text Extraction (VTE) Kiela et al. (2020), TextVQA Singh et al. (2019), Visual Dialogue (VisDial) and Commonsense VQA (CVQA); 2) Multiple-choice task, which provides instance-dependent options for each question, including several VQA tasks (Grounded VQA Zhu et al. (2016), Science QA Lu et al. (2022), IconQA Lu et al. (2021), VizWiz Gurari et al. (2018)) plus Disaster Classification (DC) Alam et al. (2023); 3) Discriminative tasks, including Visual Entailment, Visual Spatial Reasoning and Natural Language for Visual Reasoning, which can be viewed as a special variant of multiple-choice or multiple-choice with fixed options. For example, the options are restricted within “yes/no/not sure” in VE and “yes/no” in other tasks. Following the evaluation process in these papers, we report Rouge-L Lin (2004) score for generative tasks and accuracy for the other two types of tasks.
For each test task, we also determine whether the instruction tuning set includes the same task (task overlap) or if the test dataset overlaps with both the tuning and pre-training datasets (data overlap), based on the visible data sources reported in the corresponding dataset papers Xu et al. (2023b); Dai et al. (2023), as illustrated in Table 10. This could help us make an in-depth analysis. Regarding each base model, we only report the held-out tasks for fair comparison. To this end, we exclude the results of VTE and VisDial on IBLIP in Table 1.
Source | Task Name | Type | MIns | IBLIP | ||
T | D | T | D | |||
MIns | VTE | G | ||||
TextVQA | G | |||||
VisDial | G | |||||
CVQA | MC | |||||
DC | MC | |||||
Grounded VQA | MC | |||||
VE | D | |||||
VSR | D | |||||
NLVR | D | |||||
IBLIP-Bench | ScienceQA | MC | ||||
IconQA | MC | |||||
VizWiz | MC |
D.4 Evaluation Methods
We describe the evaluation formulation for all the tasks and models respectively. The tasks can be generally divided into 2 types: generation tasks and classification tasks (including multiple-choice tasks).
Generation Tasks
The evaluation for generation tasks on both models are same and pretty straightforward. We let the model generate responses under fixed hyperparameters (e.g., beam size, maximum and minimum length) and compute the ROUGE-L score with ground truths.
Classification Tasks
Evaluation for classification tasks is quite different using the two models. For OFA, we still formulate as generation tasks and the accuracy is defined as:
For IBLIP, we calculate the sequence-to-sequence CE loss from prompt to the option-appended prompt. If an instance has options, then the criterion is: the prediction is correct if the truth-appended prompt earns the minimum loss among all options,
where and is the index of truth (answer) in the given option list.
D.5 Templates for IBLIP-Bench Tasks
We provide the evaluation templates for two multiple choice IBLIP-Bench tasks (Science QA, Icon QA) in Table 11. The raw instruction can be found in Dai et al. (2023), which we show in the row “IBLIP”. The instructions in the “OFA” row are what we create from the raw instruction by modifying delimiters and special tokens to be compatible with OFA’s expression style. Here to simplify the expression, we assume there are two options, “Apple” and “Pear” and the correct answer is always “Apple”. We also list the target strings to judge whether the output is correct.
Task | Style | Template | Target |
Science QA | IBLIP | Context: {C} Question: {Q} Options: Apple||||Pear. Answer: | Apple |
OFA | {C} Question: {Q} [Options]: Apple||||Pear | Apple | |
Icon QA | IBLIP | Question: {Q} Options: Apple||||Pear. Short answer: | Apple |
OFA | {Q} [Options]: Apple||||Pear | Apple |
Appendix E Result Visualization
In this section, we visualize and compare the results tuned on MIns and MIns+ series in Table 12 and Table 13. In both cases, the MLLM either misunderstands the target of the task or fails to predict the correct answer.
![[Uncaptioned image]](x7.png)
![[Uncaptioned image]](x8.jpg)